OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10

OpenStudy (aroub):

10-1/1+2 9/3 3 y=mx+b, m=3 , (1,10) 10=3(1)+b 10-3=b 7=b y=3x-7

OpenStudy (anonymous):

This is not an option, good try though

OpenStudy (aroub):

oh can u write the options??

OpenStudy (anonymous):

3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (aroub):

well yea its the same thing but my answer is in the simplest form so the answer is 3x - y = -7

OpenStudy (anonymous):

Solve this :) What is the x-intercept and y-intercept of the graph of 4x - 8y = 32

OpenStudy (aroub):

hehehe okay!! :) -8y=-4x+32 y=1/2x-4 slope:1/2 y-intercept:-4

OpenStudy (anonymous):

ane x-intercept would be..?

OpenStudy (aroub):

oh sorry ! ok the x-intercept is 8

OpenStudy (anonymous):

Very good ! (:

OpenStudy (aroub):

hehehe!! :D

OpenStudy (anonymous):

I am posting another question in a second !

OpenStudy (aroub):

post in in a new one!